Suppose a man standing on a flat wooden platform tries to pull himself up by pulling the two ropes downwards (refer to Fig. 1). The downward force applied on each rope is F, the mass of the man is m1 and the mass of the wooden platform is m2. Is it possible for the man to accelerate upwards? If it’s possible, then find an expression for the acceleration a (take the acceleration of gravity to be g), otherwise explain why it’s not possible.


1
Expert's answer
2021-07-22T10:02:50-0400

Solution.

"\sum F=ma;"

"F+F+F+F-(m_1+m_2)g=(m_1+m_2)a;"

4F-(m_1+m_2)g=(m_1+m_2)a;

"a=\dfrac{4F-(m_1+m_2)g}{m_1+m_2};"

Answer: "a=\dfrac{4F-(m_1+m_2)g}{m_1+m_2}."